OpenStudy (anonymous):

Simplify (4y^2-13y+3)/(2y^2-5y-12)(2y^2+9y+9)/(16y^2-1)(y^2+3y-28)/(y^2-9)

OpenStudy (campbell_st):

well start by factorising the quadratics which will give \[\frac{(4y -1)(y -3)}{(2y + 3)( y - 4)} \times \frac{(2y + 3)(y + 3)}{(4y -1)(4y + 1)} \times \frac{(y + 7)(y - 3)}{(y - 3)(y+3)}\] now just cancel common factors... the distribute to get the final simplification.

OpenStudy (anonymous):

Incorrect @campbell_st \[\frac{ (4y-1)(y-3) }{ (2y+3)(y-4) }\times \frac{ (2y+3)(y+3) }{ (4y-1)(4y+1) }\times \frac{ (y+7)(y-4) }{ (y-3)(y+3) }\]

OpenStudy (anonymous):

You incorrectly factorised the numerator of the last fraction.

OpenStudy (anonymous):

\[=y+7\]